Aluminum maker Alcoa has a beta of about 1.89​, whereas Hormel Foods has a beta of 0.42.

If the expected excess return of the market portfolio is 6​%,

which of these firms has a higher equity cost of​ capital, and how much higher is​ it?

​Alcoa's equity cost of capital is? (Round to two decimal​ places.)
Answer #1

​Alcoa's equity cost of capital = 1.89 × 0.06 = 0.1134 or 11.34%

Hormel Foods equity cost of capital = 0.42 × 0.06 = 0.0252 or 2.52%

Alcoa has a higher equity cost of​ capital by 8.82%.